You and your friend live 12 miles apart. You leave home at the same time and travel toward each other. You walk at a rate of 4 miles per hour and your friend bicycles at a rate of 11 miles per hour.
a) How long after you leave home will you meet? How far from home will each of you be?
DIfference in the rate is 7 m/h. from there I

Answer by jorel1380(3719)

n(4+11)=12
15n=12
n=12/15 hours=4/5=48 minutes
You're walking at 4 mph for 4/5 hour, which is 16/5 miles. Your friend biked at 11 mph for 4/5 hours, which is 44/5 miles.